# New GPU developments in the Madgraph CUDACPP plugin: kernel splitting, helicity streams, cuBLAS color sums

### v1

@@ 3.4 Feynman diagrams as separate kernels (“ihel4”)
The final step in my developments consisted in trying to further decompose the calculate_jamps kernel, which in ihel3 computes dual amplitudes from all Feynman diagrams, into many shorter kernels. The natural way to address this, without major changes in the algorithm, consisted in encapsulating the calculation of every Feynman diagram in an individual GPU kernel. In practice, rather than launching one calculate_jamps kernel per helicity stream, which internally processes all Feynman diagrams, the new code launches many kernels diagram1, diagram2, …diagramN sequentially in each helicity stream. In both cases, the processing of Feynman diagrams must follow a predefined order, because for every diagram the algorithm knows which intermediate wavefunctions have already been computed, and which wavefunctions must still be computed. This “ihel4” version of the software is illustrated in Fig. 3.
@@
Unfortunately, this version of the software has a consistently worse performance than the previous ones, both on NVidia and AMD GPUs, as shown in Figures 4 and 5. With respect to ihel3, for instance, the ihel4 throughputs for $gg\!\rightarrow\!t\bar{t}ggg$ on an NVidia V100 GPU are a factor 3-4 worse than ihel3 at their peak for large grids, and worse by much larger factors for small grids. In addition, ihel4 crashes for very large grids, in configurations where ihel3 performs very well. The ihel4 software in vectorized C++ for SIMD CPUs also performs worse than ihel3, with degradations in throughput of the order of 20% to 50%.
One possible explanation for the worse GPU performance of ihel4 is more frequent memory access. One technical challenge in this case, in fact, is that the wavefunctions for all external and internal particles, which are a local variable wf in the calculate_wavefunctions and calculate_jamps kernels in versions ihel0–ihel3 of the code (see Table 1), must now also be stored into and retrieved from GPU global memory. In addition, as previously mentioned in a footnote, the existing implementation of helicity amplitude functions like FFV1_0 essentially constrains the wavefunction layout in GPU global memory to be an Array-Of-Structures (AOS) wf[nevt][6][2], which is suboptimal as it does not allow coalesced memory access as an SOA would do. This further increases the pressure on the memory bandwidth, which may be one cause of slowdowns.

### v2

@@ 4.1 Feynman diagrams as individual kernels (“ihel4”)
The first step in my attempt to split the calculation of all Feynman diagrams in calculate_jamps kernel into smaller GPU kernels consisted, not surprisingly, in encapsulating each diagram in an individual kernel. In practice, rather than launching one calculate_jamps kernel per helicity stream, which internally processes all Feynman diagrams, the new code launches many kernels diagram1, diagram2, …diagramN sequentially in each helicity stream. In both cases, the processing of Feynman diagrams must follow a predefined order, because for every diagram the algorithm knows which intermediate wavefunctions have already been computed, and which wavefunctions must still be computed. This “ihel4” version of the software is illustrated in Fig. 9.
@@
Unfortunately, this version of the software has a consistently worse performance than the previous ones, both on NVidia and AMD GPUs, as shown in Figures 4 and 5. With respect to ihel3, for instance, the ihel4 throughputs for $gg\!\rightarrow\!t\bar{t}ggg$ on an NVidia V100 GPU are a factor 3-4 worse than ihel3p1 at their peak for large grids, and worse by much larger factors for small grids. In addition, ihel4 crashes for very large grids, in configurations where ihel3p1 performs very well. As shown in Fig. 6, the ihel4 software in vectorized C++ for SIMD CPUs also performs worse than ihel3p1, with degradations in throughput up to 40% for $gg\!\rightarrow\!t\bar{t}ggg$.
There are at least two possible explanations for the worse GPU performance of ihel4: heavier access to GPU global memory, and kernel launch overhead. At the time of writing my previous preprint bib:preprintV1 for this article, I wrongly assumed that the latter was the most likely culprit. With hindsight and after many additional tests and developments, it is now clear that the bottleneck is instead the former. Kernel launch overhead can be reduced by adopting the CUDA graphs bib:graphsblog technology: this is what I implemented in the ihel4p1 software, described below. Reducing the impact of GPU global memory access, conversely, required a much more radical rethink of Feynman diagram splitting, described in the two following subsections: first of all, by launching kernels which include groups of many Feynman diagrams (ihel5), rather than individual diagrams; later on, by moving back to a single kernel, which however is no longer monolithic as it internally calls diagram groups as GPU device functions (ihel6).
Figure 11: Schematic representation of the CUDACPP ME engine (sigmaKin), in the last kernel splitting scenario described in this paper: (ihel6) Feynman diagram groups as individual kernels or as device functions in a single kernel.
4.1.1 Small kernels in CUDA graphs (“ihel4p1”)